in a solvent free Baeyer-Villiger reaction, the mass of the starting material or 4-tert-butylcyclohexanone is 0.498g, the mass of m-chloroperoxybenzoic acid (reagent) is 1.132g and the mass of the product y-tert-butyl-e-caprolactone is 0.297g. could you explain the limiting reagent and equivalents (show caluclations) ?
Answer #1

ANSWER:-

limiting reactant is that reactant in which less number of moles are involved.

so for calculation number of moles we use the formula n=w/m

so for 4-tert-butylcyclohexanone w = 0.498 g

molar mass of 4-tert-butylcyclohexanone 154.25 g/mol

so number of moles = 0.498/154.25

n = 0.003229 moles that is 3.22 mmol of 4-tert-butylcyclohexanone

number of moles for m-chloroperoxybenzoic acid = w/m

mass of m-chloroperoxybenzoic acid 172.57 g/mol

n = 1.132 / 172.57 = 0.00656 moles that is 6.56 mmol of m-chloroperoxybenzoic acid

so here limiting reactant is 4-tert-butylcyclohexanone because of less moles (3.22)

now determination of molar equivalents for determine the molar equivalents we should know the limiting reactant and their moles so here we know that limiting reactant is 4-tert-butylcyclohexanone (3.22 mmol).

calculate molar equivalents for each reagent, divide the moles of that reagent by the moles of the limiting reagent

molar equivalents for 4-tert-butylcyclohexanone = 3.22/3.22 = 1 equivalent

molar equivalents for m-chloroperoxybenzoic acid = 6.56mmol/3.22mmol = 2.04 equivalents.